# Optimizing post-Newtonian parameters and fixing the BMS frame for numerical-relativity waveform hybridizations

### v1

@@ VII.1 Spin-aligned systems
Figure 12 shows how we determine the end time of the matching window for spin-aligned systems. In this figure, we fix the length of the matching window as 15 orbits long, and vary the end time of the matching window. We plot $\mathcal{E}$ in both the matching window and the test window as blue and green curves, respectively. We can see that there is an optimal choice for the end time of the matching window, which is about $3000\,M$ to $8000\,M$ before merger, where the errors in the matching and test windows show a minimum. For matching windows ending earlier than this optimum, the errors are larger because of larger eccentricity, as shown in Sec. VI. The slope of the blue curve before $-5000\,M$ agrees very well with eccentricity damping, which is shown as the red dotted curve. For matching windows ending later than the optimum, we get larger errors because PN breaks down at late times.
@@
To find the minimal duration of the matching window, we fix the end time based on Figure 12, and change the start time. We also use a test window to check the results. The position of the matching window and test window are shown in the upper panel of Fig. 17, and the hybridization errors in both windows are shown in the lower panel. The solid, dashed, and dotted curves in the lower panel show the errors using NR waveforms with different numerical resolution, labeled by “Lev” as in Section VI above. We can see from the blue curves that the NR numerical error does not significantly affect the hybridization error in the matching window. But when comparing the green curves, we can see that the errors in the test window roughly converge to smaller values as we increase the resolution. This suggests that low-resolution NR simulations can accumulate numerical phase error over long periods of time.
@@
We can see from the solid green and blue curves that the optimal length of matching window should be around 5 to 20 orbits. The shorter the matching window is, the less information it includes, and the larger the chance that the small error in the matching window is spurious; for very short matching windows ($<5$ orbits) the green curves no longer converge and the solid green curve shows a large error in the test window. For longer matching windows ($>20$ orbits), the green solid curve becomes flat and the blue curves increase. This is because on the one hand, the start time of the matching window is earlier, where the NR eccentricity is larger, resulting in a larger error in the matching window, as shown by the blue curve in Fig. 12. On the other hand, a longer matching window leads to more accumulated phase error, potentially caused by both the disagreement between zero-eccentricity PN and small-but-finite-eccentricity NR, as well as insufficient NR resolution. In Fig. 18, we plot the residual corresponding to the rightmost points for Lev3 in Fig. 17. We can see that even if the matching window is long enough to be adjacent to the test window, we still cannot achieve good alignment between PN and NR outside the matching window. The error increases beyond the matching window because the PN and NR waveforms are inherently not identical. In this case of spin-aligned systems, we use a non-eccentric PN model to hybridize with NR waveforms that possess slight eccentricity. Given the inherent differences between the waveform models, it is not possible to reduce the error outside the matching window, even if we achieve minimal error within it.
@@
In Fig. 12, the blue curve before $-8000\,M$ corresponds to the LHS, the blue curve after $-3000\,M$ corresponds to the RHS, and the region between $-8000\,M$ and $-3000\,M$ is where the two effects are comparable. For cases with different but small eccentricities, we expect the error curve after $-3000\,M$ to be similar to that in Fig. 12, and the curve before $-8000\,M$ to be parallel to the blue curve in Fig. 12, but with a vertical shift. Since the curve after $-3000\,M$ is steep and the curve before $-8000\,M$ is gradual, the point where the left and right part of the curve meet with each other (which is the best location for the matching window) is not very sensitive to the eccentricity of the system.
At the current state-of-the-art NR accuracy, and for $\mathcal{E}$ between PN and NR in the test window of $\sim 10^{-4}$, the minimal choice for a matching window for spin-aligned systems would be at least $5$ orbits long and ending $3000\,M$ before merger. This minimal choice will become longer and earlier if higher-accuracy hybrids are desired, and if the accuracy of the waveforms improves. Matching windows longer than 20 orbits or ending earlier than $8000\,M$ are not beneficial given the current state of the art, but would be beneficial if the residual eccentricity of NR waveforms were decreased or if eccentric PN waveforms were used for hybridization.
@@ VII.2 Precessing systems
In Sec. VI.2, we established that the error in hybridizing precessing systems mainly results from the absence of higher-order PN spin terms. Because PN error decreases with smaller frequencies, we anticipate smaller hybridization errors at earlier times. To determine the best location of the matching window, we plot the hybridization error of system SXS:BBH:2621 in the matching window versus orbital angular velocity $\Omega_{\text{orb}}$ in Fig. 19. This system has $\chi_{p}=0.49$. Since the leading missing PN precession-related spin terms are expected to be $\propto(S_{\perp}/M^{2})^{\alpha}v^{\beta}$, where $v=(\Omega_{\text{orb}}M)^{1/3}$ is the PN expansion parameter, we can get the indices $\alpha$ and $\beta$ by fitting the points in Fig. 15 and Fig. 19, respectively. Then we can get an empirical formula $\mathcal{E}\sim 5\times 10^{8}\times(S_{\perp}/M^{2})^{3.8}\Omega_{\text{orb}}^{5}$, or
$\Omega_{\text{orb}}\lesssim 0.018\mathcal{E}^{0.2}(S_{\perp}/M^{2})^{-0.76}.$ (15)
Once a required precision $\mathcal{E}$ for a 10-orbit long window is given, we can use Eq. (15) to determine the average orbital angular velocity of the matching window. From the average orbital angular velocity one can figure out the end time of the matching window. Note that both Figs. 15 and 19 use a 10-orbit long matching window, because it is short enough to avoid being dominated by long-term accumulated phase error, and also limits significant frequency changes within the window.
@@ VII.3 Short waveforms
In Fig. 21, we display $\mathcal{E}$ for different lengths of the matching window for system SXS:BBH:2617 for this scenario. The dotted lines in the upper panel show that the error in the matching window for extrapolated waveforms is independent of the length of the matching window because the error is dominated by the incorrect gravitational-wave memory in the extrapolated waveforms. The solid lines show that as the length of the matching window becomes shorter, the error in the matching window becomes smaller. However, by comparing the blue and green curves between the upper and lower panels, we can see that smaller error in the matching window does not mean smaller error in the test window. To more intuitively show the relation between the error in the matching window and the test window, we take all the blue and green points in Fig. 21, and plot them in Fig. 22. We can see that the errors in the test window and matching window show no unambiguous trend. This non-correlation of the matching and test window errors suggests that although we can always get a small error in the matching window if the matching window is short, it is not gaining us more information about the PN parameters and BMS frame to improve the error in the test window.
@@ VIII Conclusions
However, a more informative measure of the accuracy of a hybridized waveform is the PN-NR difference in a test window that is earlier than the matching window so that PN errors are small, and that does not overlap with the matching window so that we are not actively minimizing the PN-NR difference there. In a test window that is 10 orbits long and 30 orbits earlier than the matching window, we can reduce the normalized PN-NR difference to less than $\sim 10^{-4}$ for spin-aligned systems, as shown by the solid green curve in Fig. 17. This corresponds to $\sim 10^{-2}$ absolute difference between PN and NR waveforms in this interval. Again, we note that these numbers are not directly comparable to the frequency-domain mismatches used in GW data analysis, even assuming a flat noise curve, because the test and matching windows will not generally encompass the entire detectable waveform. At best, these numbers constitute lower bounds for the full mismatch. However, they do suggest that even for spin-aligned systems the results may be inadequate for high-SNR events in future detectors such as Cosmic Explorer and LISA.
@@
Hybridization of precessing systems is much more difficult than spin-aligned systems, because the current PN waveforms for precessing systems are much less accurate than for spin-aligned systems. For precessing systems, the error in the hybrid is dominated by the lack of precession-related spin-spin terms in the PN model. Unlike the error caused by non-zero eccentricity, the error stemming from PN terms diminishes at earlier times. So there is no optimal choice of the matching window even in the entire $10^{5}\,M$ span of the NR waveforms in the SXS catalog, and it always improves the hybridization by using matching windows as long and early as possible. We provide a formula, Eq. 15, and a plot, Fig. 20, as a guide for choosing a matching window for precessing hybrids. The recommended end time for the matching window is significantly earlier for systems with large spin.
@@ Acknowledgements.
We thank Sizheng Ma, Nils Deppe, Qing Dai and Harald Pfeiffer for useful discussions. Computations for this work were preformed on the Wheeler cluster at Caltech, which is supported by the Sherman Fairchild Foundation and by Caltech, Resnick High Performance Computing (HPC) Cluster at the Caltech High Performance Computing Center, Frontera at the Texas Advanced Computing Center, and Urania HPC system at the Max Planck Computing and Data Facility. This work was supported in part by the Sherman Fairchild Foundation, by NSF Grants PHY-2207342 and OAC-2209655 at Cornell, and by NSF Grants PHY-2309211, PHY-2309231, and OAC-2209656 at Caltech. The work of L.C.S. was partially supported by NSF CAREER Award PHY-2047382 and a Sloan Foundation Research Fellowship. V.V. acknowledges support from NSF Grant No. PHY-2309301; UMass Dartmouth’s Marine and Undersea Technology (MUST) Research Program funded by the Office of Naval Research (ONR) under Grant No. N00014-23-1–2141; and the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 896869.

### v2

@@ VII.1 Spin-aligned systems
Figure 14 shows how we determine the end time of the matching window for spin-aligned systems. In this figure, we fix the length of the matching window as 15 orbits long, and vary the end time of the matching window. We plot $\mathcal{E}$ in both the matching window and the test window as blue and green curves, respectively. We can see that there is an optimal choice for the end time of the matching window, which is about $3000\,M$ to $8000\,M$ before merger, where the errors in the matching and test windows show a minimum. For matching windows ending earlier than this optimum, the errors are larger because of larger eccentricity, as shown in Sec. VI. The slope of the matching window error (blue curve) before $-5000\,M$ agrees very well with eccentricity damping, which is shown as the red dotted curve. For matching windows ending later than the optimum, we get larger errors because PN breaks down at late times.
@@
To find the minimal duration of the matching window, we fix the end time based on Figure 14, and change the start time. We also use a test window to check the results. The position of the matching window and test window are shown in the upper panel of Fig. 19, and the hybridization errors in both windows are shown in the lower panel. The solid, dashed, and dotted curves in the lower panel show the errors using NR waveforms with different numerical resolution, labeled by “Lev” as in Section VI above. We can see from the matching window errors (blue curves) that the NR numerical error does not significantly affect the hybridization error in the matching window. But when comparing the test window errors (green curves), we can see that they roughly converge to smaller values as we increase the resolution. This suggests that low-resolution NR simulations can accumulate numerical phase error over long periods of time. But note that the NR dephasing between the two highest Levs, $\Delta\phi$, does not directly reflect the inaccuracy of the NR waveforms, as the NR parameters within the matching window for different Levs can deviate by $10^{-2}$, given their evolution over nearly a hundred orbits prior to the matching window. Consequently, the optimized PN parameters for various NR Levs also differ by $10^{-2}$, allowing the PN waveform to agree better with their respective NR waveforms and partially compensating for the phase error. Assessing the adequacy of the current NR resolution for hybridization purposes requires performing the entire hybridization process using various NR Levs, as shown in Fig. 19. However, the conclusion depends on the precision and frequency band needed for these observations (specifically, the frequency band of the test window). We leave this for future work.
@@
We can see from the Lev3 errors (solid green and blue curves) that the optimal length of matching window should be around 5 to 20 orbits. The shorter the matching window is, the less information it includes, and the larger the chance that the small error in the matching window is spurious; for very short matching windows ($<5$ orbits) the test window errors (green curves) no longer converge and the Lev3 test window error (solid green curve) increases significantly. For longer matching windows ($>20$ orbits), the Lev3 test window error (green solid curve) becomes flat and the matching window errors (blue curves) increase. This is because on the one hand, the start time of the matching window is earlier, where the NR eccentricity is larger, resulting in a larger error in the matching window, as shown by the blue curve in Fig. 14. On the other hand, a longer matching window leads to more accumulated phase error, potentially caused by both the disagreement between zero-eccentricity PN and small-but-finite-eccentricity NR, as well as insufficient NR resolution. In Fig. 20, we plot the residual corresponding to the rightmost points for Lev3 in Fig. 19. We can see that even if the matching window is long enough to be adjacent to the test window, we still cannot achieve good alignment between PN and NR outside the matching window. The error increases beyond the matching window because the PN and NR waveforms are inherently not identical. In this case of spin-aligned systems, we use a non-eccentric PN model to hybridize with NR waveforms that possess slight eccentricity. Given the inherent differences between the waveform models, it is not possible to reduce the error outside the matching window, even if we achieve minimal error within it.
@@
In Fig. 14, the matching window error (blue curve) before $-8000\,M$ corresponds to the left-hand side, the matching window error after $-3000\,M$ corresponds to the RHS, and the region between $-8000\,M$ and $-3000\,M$ is where the two effects are comparable. For cases with different but small eccentricities, we expect the error curve after $-3000\,M$ to be similar to that in Fig. 14, and the curve before $-8000\,M$ to be parallel to the blue curve in Fig. 14, but with a vertical shift. Since the curve after $-3000\,M$ is steep and the curve before $-8000\,M$ is gradual, the point where the left and right part of the curve meet with each other (which is the best location for the matching window) is not very sensitive to the eccentricity of the system.
At the current state-of-the-art NR accuracy, and for $\mathcal{E}$ between PN and NR in the test window of $\sim 10^{-4}$, the minimal choice for a matching window for spin-aligned systems would be at least $5$ orbits long and ending $3000\,M$ before merger. This minimal choice will become longer and earlier if higher-accuracy hybrids are desired, and if the accuracy of the waveforms improves. Matching windows longer than 20 orbits or ending earlier than $8000\,M$ are not beneficial given the current state of the art, but would be beneficial if the residual eccentricity of NR waveforms were decreased or if eccentric PN waveforms were used for hybridization and the potential degeneracy between eccentricity and supertranslations are properly handled.
@@ VII.2 Precessing systems
In Sec. VI.2, we established that the error in hybridizing precessing systems mainly results from the absence of higher-order PN spin terms. Because PN error decreases with smaller frequencies, we anticipate smaller hybridization errors at earlier times. To determine the best location of the matching window, we plot the hybridization error of system SXS:BBH:2621 in the matching window versus orbital angular velocity $\Omega_{\text{orb}}$ in Fig. 21. This system has $\chi_{p}=0.49$. Since the leading missing PN precession-related spin terms are expected to be $\propto(S_{\perp}/M^{2})^{\alpha}v^{\beta}$, where $v=(\Omega_{\text{orb}}M)^{1/3}$ is the PN expansion parameter, we can get the indices $\alpha$ and $\beta$ by fitting the points in Fig. 17 and Fig. 21, respectively. Then we can get an empirical formula $\mathcal{E}\sim 3.6\times 10^{8}\times(S_{\perp}/M^{2})^{4.1}\Omega_{\text{orb}}^{5}$, or
$\Omega_{\text{orb}}\lesssim 0.012\mathcal{E}^{0.2}(S_{\perp}/M^{2})^{-0.82}.$ (15)
Once a required precision $\mathcal{E}$ for a 10-orbit long window is given, we can use Eq. (15) to determine the average orbital angular velocity of the matching window. From the average orbital angular velocity one can figure out the end time of the matching window. Note that both Figs. 17 and 21 use a 10-orbit long matching window, because it is short enough to avoid being dominated by long-term accumulated phase error, and also limits significant frequency changes within the window.
@@ VII.3 Short waveforms
In Fig. 23, we display $\mathcal{E}$ for different lengths of the matching window for system SXS:BBH:2617 for this scenario. The matching window errors obtained using extrapolated waveforms (dotted lines) in the upper panel are independent of the length of the matching window because the error is dominated by the incorrect gravitational-wave memory in the extrapolated waveforms. The matching window errors obtained using CCE waveforms (solid lines) show that as the length of the matching window becomes shorter, the error in the matching window becomes smaller. However, by comparing the errors in the matching window and test window (upper and lower panels), we can see that smaller error in the matching window does not mean smaller error in the test window. To more intuitively show the relation between the error in the matching window and the test window, we gather all the data points from Fig. 23, and plot them again in Fig. 24. We can see that the errors in the test window and matching window show no unambiguous trend. This non-correlation of the matching and test window errors suggests that although we can always get a small error in the matching window if the matching window is short, it is not gaining us more information about the PN parameters and BMS frame to improve the error in the test window.
@@ VIII Conclusions
However, a more informative measure of the accuracy of a hybridized waveform is the PN-NR difference in a test window that is earlier than the matching window so that PN errors are small, and that does not overlap with the matching window so that we are not actively minimizing the PN-NR difference there. In a test window that is 10 orbits long and 30 orbits earlier than the matching window, we can reduce the normalized $L^{2}$ norm of PN-NR difference to less than $\sim 10^{-4}$ for spin-aligned systems, as shown by the solid green curve in Fig. 19. This corresponds to $\sim 10^{-2}$ residual between PN and NR waveforms in this interval. Again, we note that these numbers are not directly comparable to the frequency-domain mismatches used in GW data analysis, even assuming a flat noise curve, because the test and matching windows will not generally encompass the entire detectable waveform. At best, these numbers constitute lower bounds for the full mismatch. However, they do suggest that even for spin-aligned systems the results may be inadequate for high-SNR events in future detectors such as Cosmic Explorer and LISA.
@@
Hybridization of precessing systems is much more difficult than spin-aligned systems, because the current PN waveforms for precessing systems are much less accurate than for spin-aligned systems. For precessing systems, the error in the hybrid is likely to be dominated by the lack of precession-related spin-spin terms in the PN model. Unlike the error caused by non-zero eccentricity, the error stemming from PN terms diminishes at earlier times. So there is no optimal choice of the matching window even in the entire $10^{5}\,M$ span of the NR waveforms in the SXS catalog, and it always improves the hybridization by using matching windows as long and early as possible. We provide a formula, Eq. 15, and a plot, Fig. 22, as a guide for choosing a matching window for precessing hybrids. The recommended end time for the matching window is significantly earlier for systems with large spin.
It is possible that using the effective one-body (EOB) formalism [80] instead of PN could significantly affect the position and length of the matching window and decrease the requirement for long NR waveforms. However, to do so requires that one have $\Psi_{M}$ in the EOB formalism to be able to fix the BMS freedoms, which is not currently the case. Furthermore, current EOB waveforms are calibrated using the extrapolated NR waveforms, which lack memory effects. We plan to explore how the hybridization procedure changes when using EOB waveforms trained on CCE waveforms in future work.
@@ Acknowledgements.
We thank Sizheng Ma, Nils Deppe, Qing Dai, Harald Pfeiffer and Aditya Vijaykumar for useful discussions. Computations for this work were preformed on the Wheeler cluster at Caltech, which is supported by the Sherman Fairchild Foundation and by Caltech, Resnick High Performance Computing (HPC) Cluster at the Caltech High Performance Computing Center, Frontera at the Texas Advanced Computing Center, and Urania HPC system at the Max Planck Computing and Data Facility. This work was supported in part by the Sherman Fairchild Foundation, by NSF Grants PHY-2207342 and OAC-2209655 at Cornell, and by NSF Grants PHY-2309211, PHY-2309231, and OAC-2209656 at Caltech. The work of L.C.S. was partially supported by NSF CAREER Award PHY-2047382 and a Sloan Foundation Research Fellowship. V.V. acknowledges support from NSF Grant No. PHY-2309301; UMass Dartmouth’s Marine and Undersea Technology (MUST) Research Program funded by the Office of Naval Research (ONR) under Grant No. N00014-23-1–2141; and the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No. 896869.